# FedRSU: Federated Learning for Scene Flow Estimation on Roadside Units

### v1

@@ Abstract
Roadside unit (RSU) can significantly improve the safety and robustness of autonomous vehicles through Vehicle-to-Everything (V2X) communication. Currently, the usage of a single RSU mainly focuses on real-time inference and V2X collaboration, while neglecting the potential value of the high-quality data collected by RSU sensors. Integrating the vast amounts of data from numerous RSUs can provide a rich source of data for model training. However, the absence of ground truth annotations and the difficulty of transmitting enormous volumes of data are two inevitable barriers to fully exploiting this hidden value. In this paper, we introduce FedRSU, an innovative federated learning framework for self-supervised scene flow estimation. In FedRSU, we present a recurrent self-supervision training paradigm, where for each RSU, the scene flow prediction of points at every timestamp can be supervised by its subsequent future multi-modality observation. Another key component of FedRSU is federated learning, where multiple devices collaboratively train an ML model while keeping the training data local and private. With the power of the recurrent self-supervised learning paradigm, FL is able to leverage innumerable underutilized data from RSU. To verify the FedRSU framework, we construct a large-scale multi-modality dataset RSU-SF. The dataset consists of 17 RSU clients, covering various scenarios, modalities, and sensor settings. Based on RSU-SF, we show that FedRSU can greatly improve model performance in ITS and provide a comprehensive benchmark under diverse FL scenarios. To the best of our knowledge, we provide the first real-world LiDAR-camera multi-modal dataset and benchmark for the FL community.
@@ I Introduction
Roadside units (RSUs) are an important constituent of intelligent transportation systems (ITS). They are often installed along roadsides, intersections, and other transportation scenarios and outfitted with various sensors and communication devices [1, 2, 3]. With the rapid evolution of vehicle-to-infrastructure communication technologies [4], RSUs are able to offer significant added value for intelligent autonomous vehicles. RSU perception can expand the field of view of autonomous vehicles, mitigate occlusion issues, and provide more accurate perception results for crucial categories such as pedestrians and cyclists [5, 6, 7]. Additionally, RSUs can serve as the central node of vehicle-to-everything (V2X) communication, consolidating information from multiple autonomous vehicles and offering the potential to deploy more powerful computing devices for local training missions [8, 9, 3].
At present, the RSU perception system only employs fixed machine-learning models for real-time situation awareness. Generally, these models are pre-trained on limited labeled datasets. Despite the fact that RSUs are equipped with high-resolution sensors capable of providing stable and precise scene observations, the continuously collected streaming data has not been effectively utilized to adapt the model to the current perceptual scene or to improve the model’s perceptual capability over time. From the model training perspective, a substantial portion of the streaming data remains underutilized.
@@
Two formidable challenges need to be addressed in order to leverage the infinite streaming data from RSU sensors to facilitate model training. (1) The data labeling process is exceedingly arduous for transportation data. For RSU perception, the annotation of traffic data is crucial and advantageous for model training. However, the annotation process is always arduous and time-consuming, requiring extensive data storage, transmission, and manual labeling. The creation of annotated datasets [10, 11, 12, 13, 14] is always a laborious and expensive endeavor. It is unrealistic to annotate all the data collected by RSUs to continuously refine the perception model. (2) Integrating data or models from multiple RSU edges is not trivial. In the intelligent transportation system, it is challenging to gather a massive amount of unlabeled data from various sources and train a universal model, as successfully achieved in computer vision [15] or natural language processing [16]. The traffic scene data collected by RSUs may comprise a lot of sensitive information and cannot be shared publicly, such as camera data that contains sensitive facial features and vehicle license plate information. Moreover, the sheer amount of data makes it impossible to aggregate for centralized training due to communication bandwidth and storage limitations.
In our design, we designate scene flow estimation as the principal task for each RSU. Scene flow [17], which is the 3D equivalent of optical flow [18] and describes the motion vector of points in 3D space. Reasoning about motion and forecasting the future of a dynamic scene is particularly critical. Scene flow estimation can be a crucial task that supports various downstream tasks, including motion segmentation [19], instance segmentation [20], object detection [21], motion prediction [22, 23], trajectory prediction [24], and more. Also, scene flow is a commonly used representation format in self-supervised learning methods [25, 26, 27].
To address the challenge of traffic data annotation, our core idea is to use a recurrent self-supervision paradigm on every RSU client for scene flow estimation learning, as is shown in Fig. 2. In transportation scenarios, one popular strategy of self-supervised learning is to obtain model guidance from sequential data [28, 29, 21]. With the constant stream of data, real-time scene flow prediction results of the deployed model can be supervised by subsequent future observations from RSU sensors. For each frame of the point cloud, its corresponding scene flow can be directly supervised by the succeeding frame of raw point cloud data. In this way, the recurrent self-supervised learning paradigm consumes real-time sequential data, allowing for continuous learning and improvement. Additionally, given that RSUs are typically equipped with multi-modal sensors, we propose a multi-modal self-supervised learning method that leverages consecutive images to enhance the learning of point cloud scene flow.
@@
Integrating both solutions mentioned above, in this paper, we propose FedRSU, an FL framework for self-supervised scene flow estimation in the RSU scenarios. The FedRSU framework consists of numerous RSU clients and a centralized server. On each RSU client, a scene flow estimation model is deployed and the model is continuously fine-tuned in a self-supervised manner. We propose a multi-model self-supervised method that trains the scene flow estimation model by leveraging information from both point cloud and image, where the sequential image data is applied to provide fine-grained supervision on the dynamic status of points. Besides all RSU clients, a central cloud server coordinates models among all clients. All trained local models on RSU clients are uploaded, aggregated, and combined into a single global model. The global model encapsulates group knowledge and is subsequently distributed back to each client for further refinement. Through multiple rounds of client-server interactions, the resulting global model acquires an enhanced ability to comprehend diverse real-world scenes, a feat that is challenging for a single RSU to accomplish due to its limited perceptible area. Consequently, our proposed FedRSU can enable more accurate and reliable perception capabilities in intelligent transportation systems, enhancing the effectiveness and security of the system.
@@
Previous attempts applying federated learning (FL) to ITS have covered various tasks and applications, including perception tasks (e.g., object detection [33, 34, 35], semantic segmentation [36, 37], and bird’s eye view segmentation [38]), prediction tasks (e.g., steering wheel angle prediction [39, 40] and trajectory prediction [41]), and motion control [42] or driver monitoring applications [43, 44]. However, previous efforts have serious problems in their designs and frameworks, whereas our proposed FedRSU boasts three distinct advantages over them. 1) The scene flow estimation task is critical and helpful for 3D scene understanding, which directly benefits RSU perception and overall ITS development. 2) We use RSUs as FL clients instead of onboard devices. Onboard devices have limited computing capability and cannot be employed for model training while carrying out real-time inference tasks in complex and dynamic traffic scenarios. 3) Previous works introduce the assumption of client-side ground truth labels, while raw sensor data generated by traffic participants are unable to be annotated due to privacy and data transmission limitations. Consequently, FL with supervised tasks cannot be practically deployed in a real-world ITS.
@@
Previous RSU datasets [10, 11, 12, 13, 14] only contain data from a single scenario and provide 3D bounding box annotation for scene perception. To validate our scene flow FL framework FedRSU, we propose a large-scale RSU scene flow dataset, RSU-SF. The RSU-SF dataset comprises a total of 31,311 samples collected from 17 RSUs, with each RSU acting as a client in FL. The RSU scenes are derived from three RSU datasets [10, 11, 12] and a self-collected dataset, providing a diverse range of scenarios and data heterogeneity. Data heterogeneity in the RSU-SF dataset arises from various aspects, including the diverse scenarios in which the RSUs are deployed, the different classes and densities of objects in these scenarios, and the varied modalities, sensor parameters, and sensor settings of the RSUs.
@@ II-A Roadside Units in Intelligent Transportation Systems
The domain of RSU perception in ITS requires specialized data and methods, given that the data format for RSU perception differs significantly from that of general autonomous driving datasets [58, 59]. The sensors in autonomous driving datasets are typically mounted relatively close to the ground, whereas RSUs provide an aerial perspective with a broader field of view and fewer occlusions. Consequently, numerous datasets [13, 11, 12, 14] have been proposed for this field, and many perception methods [60, 61, 62, 63] have been developed to cater to these specific scenarios. [60] predicts the height of the 3D object to enhance monocular 3d detection. [62] focuses on the camera-lidar fusion problem for RSU. Besides, a large number of studies have explored the collaboration setting [10, 64, 65] and algorithms [66, 6] between RSUs and autonomous vehicles.
@@ III-B Scene Flow Estimation via Multi-modal Recurrent Self-Supervised Learning
With this predicted point cloud, one basic loss candidate is single-modal Chamfer loss, which enforces the source point cloud to move toward the target point cloud based on point cloud only. However, despite point cloud can offer rich 3D structural information, it has limitations such as lack of color information, ambiguities in object boundaries, noise and artifacts, which can hinder the accurate prediction of scene flow.
@@
which assigns larger punishment for those points that are more likely to be static. In this way, the optical flow (image data) can assist the self-supervised learning on point cloud data providing more fine-grained supervision guidance.
Besides, we also apply a smoothness regularization to avoid being stucked in local minima [27, 26] and preserve local neighbor structures. Using $\ell_{2}$ distance, the smoothness regularization loss is defined as:
$\mathcal{L}_{reg}(\bm{\theta})=\sum_{p_{a}\in\mathbf{P}}\frac{1}{\left|\mathcal{N}(p_{a})\right|}\sum_{p_{k}\in\mathcal{N}(p_{a})}\left\|f_{a}-f_{k}\right\|_{2}^{2}.$
@@ III-C Federated Learning System
Addressing this, we propose to a new collaborative learning system FedRSU, which enables collaborative training of scene flow estimation model among multiple RSUs. Through FedRSU, the final model is essentially the union of knowledge captured by multiple RSUs, breaking the perceptual limits of each individual RSU, thus facilitating more accurate and robust estimation. Following the conventional procedures of federated learning [46], the FedRSU consists of four iterative step for each communication round, namely, global model broadcasting, local model training, local model uploading, and global model updating. The overview of the FedRSU system is shown in Fig. 3.
Global model broadcasting. In this step, the global model is broadcast to clients to serve for local model initialization. Specifically, at the beginning of each round $t$, the server broadcasts the global model $\bm{\theta}^{t}$ to each available client $k\in\mathcal{S}^{t}$, where $\mathcal{S}^{t}$ consists of index of all available clients at round $t$. Then, each client $k$ uses this global model to initialize its local model $\bm{\theta}_{k}^{(t,0)}:=\bm{\theta}^{t}$, where $\bm{\theta}_{k}^{(t,0)}$ denotes the local model at $t$-th round and 00-th training iteration.
Local model training. In this step, each client uses its local dataset to train a local model for scene flow estimation, which is guided by the self-supervised loss terms illustrated in Section III-B. Specifically, starting from initial local model $\bm{\theta}_{k}^{(t,0)}$, client $k$ will conduct multiple iterations of SGD updates based on the local dataset $\mathcal{D}_{k}$. At each iteration $r$, the local model update is represented as:
@@
Global model updating. In this step, the server updates global model by aggregating local models, which is subsequently broadcast to available clients for the next round. Specifically, the global model is weighted updated as:
@@ IV-A Dataset Collection
Scene Flow Label Generation. To validate the test the models, ground truth scene flow labels are necessary to compute and compare metrics. In RSU-SF, these labels are derived from the 3D bounding boxes and tracking labels of scene participants. The method used to generate ground truth is similar to that employed by previous methods [19]. Specifically, we extract the rigid transformation of an object from its consecutive bounding box annotations, and then the flow of points within the bounding box can be directly calculated according to the rigid transformation.
@@
Base Datasets In intelligent transportation systems, data collected from RSU sensors may differ in various aspects, such as traffic scenes, sensor modality, sensor deployment height and angle, and sensor device models and resolutions. To demonstrate the practical impact and effectiveness of our framework in real-world scenarios, the RSU-SF dataset can greatly reflect the diversity of RSUs. We select three base datasets for RSU perception [10, 11, 12] and collect another real-world RSU dataset. In total, RSU-SF consists of sensor data collected from 17 RSUs.
DAIR-V2X [10] is a large-scale multi-modality dataset for vehicle-to-infrastructure collaborative perception. The dataset comprises point cloud and image data collected from RSUs located at 7 different intersections. Due to the lack of accuracy in the original dataset labeling, we have relabeled all objects’ 3D bounding boxes and tracking labels for the infrastructure-side data. LUMPI [11] is collected using five LiDAR devices deployed at different locations around a single intersection. We consider them as separate RSUs due to their significantly varying perception ranges and resolutions. As the annotations of LUMPI are not yet released, we only use it for training. Specifically, we selected 2,400 frames for training from each RSU. IPS300+ [12] is a multimodal dataset designed for roadside perception tasks in large-scale urban intersections. The dataset was collected using two intersection perception units equipped with lidar sensors and cameras. Its data is characterized by a high density of objects in the scene, resulting in a high label density and scene complexity. We selected around 3000 frames from the original dataset, out of which 600 frames with complete labels were manually chosen for generating validation and testing pairs.
@@
TABLE I: RSU-SF setup. RSU-SF is constructed from multiple sources, which cover diverse scenarios, devices, angles, heights, and ranges.
Client Index 0-6 7 8 9 10-11 12-13 14-16
Dataset DAIR-V2X [10] LUMPI [11] IPS300+ [12] CAMPUS
Scenario Crossroads Cusy Intersection Busy Intersection Campus Roadside
Modality Image / PC PC PC PC PC PC Image / PC
With GT w w/o w/o w/o w/o w w
@@
Device Jaguar HDL64 Pandar64 PandarQT VLP16 Ruby-Lite Leishen C32
Horizontal FOV (°) 100 360 360 360 360 360 100
Vertical FOV (°) 40 26.5 40 104.2 30 40 31
Channels (beam) 300 64 64 64 16 80 32
Rate (Hz) 10 10 10 10 10 10 10
Range (m) 280 120 200 20 100 230 100
Range Accuracy (cm) 3 2 2.5 3 3 3 3
Installation Height (m) $\sim$6 $\sim$5 $\sim$5 $\sim$5 $\sim$5 5.5 $\sim$6
Installation Angle(°) 30 0 0 0 0 0 30

### v2

@@ Abstract
Roadside unit (RSU) can significantly improve the safety and robustness of autonomous vehicles through Vehicle-to-Everything (V2X) communication. Currently, the usage of a single RSU mainly focuses on real-time inference and V2X collaboration, while neglecting the potential value of the high-quality data collected by RSU sensors. Integrating the vast amounts of data from numerous RSUs can provide a rich source of data for model training. However, the absence of ground truth annotations and the difficulty of transmitting enormous volumes of data are two inevitable barriers to fully exploiting this hidden value. In this paper, we introduce FedRSU, an innovative federated learning framework for self-supervised scene flow estimation. In FedRSU, we present a recurrent self-supervision training paradigm, where for each RSU, the scene flow prediction of points at every timestamp can be supervised by its subsequent future multi-modality observation. Another key component of FedRSU is federated learning, where multiple devices collaboratively train an ML model while keeping the training data local and private. With the power of the recurrent self-supervised learning paradigm, FL is able to leverage innumerable underutilized data from RSU. To verify the FedRSU framework, we construct a large-scale multi-modality dataset RSU-SF. The dataset consists of 17 RSU clients and an additional 4 vehicle clients, covering various scenarios, modalities, and sensor settings. Based on RSU-SF, we show that FedRSU can greatly improve model performance in ITS and provide a comprehensive benchmark under diverse FL scenarios. To the best of our knowledge, we provide the first real-world LiDAR-camera multi-modal dataset and benchmark for the FL community. Code and dataset are available at https://github.com/wwh0411/FedRSU
@@ I Introduction
Roadside units (RSUs) are an important constituent of intelligent transportation systems (ITS). They are often installed along roadsides, intersections, and other transportation scenarios and outfitted with various sensors and communication devices [1, 2, 3]. With the rapid evolution of vehicle-to-infrastructure communication technologies [4], RSUs are able to offer significant added value for intelligent autonomous vehicles [5, 6, 7]. At present, the RSU perception system only employs fixed machine-learning models for real-time situation awareness. Generally, these models are pre-trained on limited labeled datasets. Despite the fact that RSUs are equipped with high-resolution sensors capable of providing stable and precise scene observations, the continuously collected streaming data has not been effectively utilized to adapt the model to the current perceptual scene or to improve the model’s perceptual capability over time. From the model training perspective, a substantial portion of the streaming data remains underutilized.
@@
Two formidable challenges need to be addressed in order to leverage the infinite streaming data from RSU sensors to facilitate model training. (1) Traffic data annotation. The data labeling process is exceedingly arduous for transportation data. For RSU perception, the annotation of traffic data is crucial and advantageous for model training. However, the annotation process is always arduous and time-consuming, requiring extensive data storage, transmission, and manual labeling. The creation of annotated datasets [8, 9, 10, 11, 12] is always a laborious and expensive endeavor. It is unrealistic to annotate all the data collected by RSUs to continuously refine the perception model. (2)Data integration. Integrating data or models from multiple RSU edges is not trivial. In the intelligent transportation system, it is challenging to gather a massive amount of unlabeled data from various sources and train a universal model, as successfully achieved in computer vision [13] or natural language processing [14]. The traffic scene data collected by RSUs may comprise a lot of sensitive information and cannot be shared publicly, such as camera data that contains sensitive facial features and vehicle license plate information. Moreover, the sheer amount of data makes it impossible to aggregate for centralized training due to communication bandwidth and storage limitations.
In our design, we designate scene flow estimation as the principal task for each RSU. Scene flow [15], which is the 3D equivalent of optical flow [16] and describes the motion vector of points in 3D space. Reasoning about motion and forecasting the future of a dynamic scene is particularly critical. Scene flow estimation can be a crucial task that supports various downstream tasks, including motion segmentation [17], object detection [18], motion prediction [19, 20], and more. Also, scene flow is a commonly used representation format in self-supervised learning methods [21, 22, 23].
To address the challenge of traffic data annotation, our core idea is to use a recurrent self-supervision paradigm on every RSU client for scene flow estimation learning, as is shown in Fig. 2. In transportation scenarios, one popular strategy of self-supervised learning is to obtain model guidance from sequential data [24, 25, 18]. With the constant stream of data, real-time scene flow prediction results of the deployed model can be supervised by subsequent future observations from RSU sensors. In this way, the recurrent self-supervised learning paradigm consumes real-time sequential data, allowing for continuous learning and improvement.
@@
Integrating both solutions mentioned above, in this paper, we propose FedRSU, an FL framework for self-supervised scene flow estimation in the RSU scenarios. The FedRSU framework consists of numerous RSU clients and a centralized server. On each RSU client, a scene flow estimation model is deployed and the model is continuously fine-tuned in a self-supervised manner. Additionally, given that RSUs are typically equipped with multi-modal sensors, we propose a multi-model self-supervised method that trains the scene flow estimation model by leveraging information from both point cloud and image. Besides all RSU clients, a central cloud server coordinates models among all clients. All trained local models on RSU clients are uploaded, aggregated, and combined into a single global model. The global model encapsulates group knowledge and is subsequently distributed back to each client for further refinement. Through multiple rounds of client-server interactions, the resulting global model acquires an enhanced ability to comprehend diverse real-world scenes, a feat that is challenging for a single RSU to accomplish due to its limited perceptible area. Consequently, our proposed FedRSU can enable more accurate and reliable perception capabilities in intelligent transportation systems, enhancing the effectiveness and security of the system.
@@
Previous attempts applying federated learning (FL) to ITS have covered various tasks and applications, including perception tasks (e.g., object detection [29, 30, 31], semantic segmentation [32, 33], and bird’s eye view segmentation [34]), prediction tasks (e.g., steering wheel angle prediction [35, 36] and trajectory prediction [37]), and motion control [38] or driver monitoring applications [39, 40]. However, previous efforts have serious problems in their designs and frameworks, whereas our proposed FedRSU boasts three distinct advantages over them. 1) The scene flow estimation task is critical and helpful for 3D scene understanding, which directly benefits RSU perception and overall ITS development. 2) We use RSUs as FL clients instead of onboard devices. Onboard devices have limited computing capability and cannot be employed for model training while carrying out real-time inference tasks in complex and dynamic traffic scenarios. 3) Previous works introduce the assumption of client-side ground truth labels, while raw sensor data generated by traffic participants are unable to be annotated due to privacy and data transmission limitations. Consequently, FL with supervised tasks cannot be practically deployed in a real-world ITS. It is important to note that FedRSU features a multi-modality nature. Our proposed dataset includes data from both LiDAR and camera sensors. Additionally, we propose a method for self-supervised learning using multi-modality data.
@@
Previous RSU datasets [8, 9, 10, 11, 12] only contain data from a single scenario and provide 3D bounding box annotation for scene perception. To validate our scene flow FL framework FedRSU, we propose a large-scale RSU scene flow dataset, RSU-SF. The RSU-SF dataset comprises a total of 31,311 samples collected from 17 RSUs, with each RSU acting as a client in FL. The RSU scenes are derived from three RSU datasets [8, 9, 10] and a self-collected dataset, providing a diverse range of scenarios and data heterogeneity. To verify and discuss whether the sensor data from vehicles connected to the RSU can effectively participate in the FedRSU system, we selected vehicle-edge data from four different intersections in the DAIR-V2X [8] dataset as 4 vehicle clients. Data heterogeneity in the RSU-SF dataset arises from various aspects, including the diverse scenarios in which the RSUs are deployed, the different classes and densities of objects in these scenarios, and the varied modalities, sensor parameters, and sensor settings of the RSUs.
@@ II-A Roadside Units in Intelligent Transportation Systems
The domain of RSU perception in ITS requires specialized data and methods, given that the data format for RSU perception differs significantly from that of general autonomous driving datasets [54, 55]. The sensors in autonomous driving datasets are typically mounted relatively close to the ground, whereas RSUs provide an aerial perspective with a broader field of view and fewer occlusions. Consequently, numerous datasets [11, 9, 10, 12] have been proposed for this field, and many perception methods [56, 57, 58] have been developed to cater to these specific scenarios. [56] predicts the height of the 3D object to enhance monocular 3d detection. [58] focuses on the camera-lidar fusion problem for RSU. Besides, a large number of studies have explored the collaboration setting [8, 59, 60] and algorithms [61, 6] between RSUs and autonomous vehicles.
@@ III-B Scene Flow Estimation via Multi-modal Recurrent Self-Supervised Learning
With this predicted point cloud, one basic loss candidate is Chamfer loss, which enforces the source point cloud to move toward the target point cloud based on point cloud only. However, despite point cloud can offer rich 3D structural information, it has limitations such as lack of color information, ambiguities in object boundaries, noise and artifacts, which can hinder the accurate prediction of scene flow.
@@
which assigns larger punishments for those points that are more likely to be static. In this way, the optical flow (image data) can assist the self-supervised learning on point cloud data providing more fine-grained supervision guidance.
Besides, we also apply a smoothness regularization to avoid being stuck in local minima [23, 22] and preserve local neighbor structures. Using $\ell_{2}$ distance, the smoothness regularization loss is defined as:
$\mathcal{L}_{reg}(\bm{\theta})=\sum_{p_{a}\in\mathbf{P}}\frac{1}{\left|\mathcal{N}(p_{a})\right|}\sum_{p_{k}\in\mathcal{N}(p_{a})}\left\|f_{a}-f_{k}\right\|_{2}^{2}.$ (2)
@@ III-C Federated Learning System
Addressing this, we propose to a new collaborative learning system FedRSU, which enables collaborative training of scene flow estimation model among multiple RSUs. Through FedRSU, the final model is essentially the union of knowledge captured by multiple RSUs, breaking the perceptual limits of each individual RSU, thus facilitating more accurate and robust estimation. Following the conventional procedures of federated learning [42], the FedRSU consists of four iterative steps for each communication round, namely, global model broadcasting, local model training, local model uploading, and global model updating. The overview of the FedRSU system is shown in Fig. 3.
Global model broadcasting. In this step, the global model is broadcast to clients to serve for local model initialization. Specifically, at the beginning of each round $t$, the server broadcasts the global model $\bm{\theta}^{t}$ to each available client $k\in\mathcal{S}^{t}$, where $\mathcal{S}^{t}$ consists of an index of all available clients at round $t$. Then, each client $k$ uses this global model to initialize its local model $\bm{\theta}_{k}^{(t,0)}:=\bm{\theta}^{t}$, where $\bm{\theta}_{k}^{(t,0)}$ denotes the local model at $t$-th round and 00-th training iteration.
Local model training. In this step, each client uses its local dataset to train a local model for scene flow estimation, which is guided by the self-supervised loss terms illustrated in Section III-B. Specifically, starting from the initial local model $\bm{\theta}_{k}^{(t,0)}$, client $k$ will conduct multiple iterations of SGD updates based on the local dataset $\mathcal{D}_{k}$. At each iteration $r$, the local model update is represented as:
@@
Global model updating. In this step, the server updates global model by aggregating local models, which is subsequently broadcast to available clients for the next round. Specifically, the global model is weighted and updated as:
@@ IV-A Dataset Collection
Scene flow label generation. To validate and test the models, ground truth scene flow labels are necessary to compute and compare metrics. In RSU-SF, these labels are derived from the 3D bounding boxes and tracking labels of scene participants. The method used to generate ground truth is similar to that employed by previous methods [17]. Specifically, we extract the rigid transformation of an object from its consecutive bounding box annotations, and then the flow of points within the bounding box can be directly calculated according to the rigid transformation.
@@
Base datasets In intelligent transportation systems, data collected from RSU sensors may differ in various aspects, such as traffic scenes, sensor modality, sensor deployment height and angle, and sensor device models and resolutions. To demonstrate the practical impact and effectiveness of our framework in real-world scenarios, the RSU-SF dataset can greatly reflect the diversity of RSUs. We select three base datasets for RSU perception [8, 9, 10] and collect another real-world RSU dataset. In total, RSU-SF consists of sensor data collected from 17 RSUs, comprising 21 clients in our dataset. Among these, data of 4 clients originate from vehicle sensors connected to RSUs.
DAIR-V2X [8] is a large-scale multi-modality dataset for vehicle-to-infrastructure collaborative perception. The dataset comprises point cloud and image data collected from RSUs and vehicles connected to the RSU located at 7 different intersections. Due to the lack of accuracy in the original dataset labeling, we have relabeled all objects’ 3D bounding boxes and tracking labels for the infrastructure-side data. LUMPI [9] is collected using five LiDAR devices deployed at different locations around a single intersection. We consider them as separate RSUs due to their significantly varying perception ranges and resolutions. As the annotations of LUMPI are not yet released, we only use it for training. Specifically, we selected 2,400 frames for training from each RSU. IPS300+ [10] is a multimodal dataset designed for roadside perception tasks in large-scale urban intersections. The dataset was collected using two intersection perception units equipped with lidar sensors and cameras. Its data is characterized by a high density of objects in the scene, resulting in a high label density and scene complexity. We selected around 3000 frames from the original dataset, out of which 600 frames with complete labels were manually chosen for generating validation and testing pairs.
@@
TABLE I: RSU-SF setup. RSU-SF is constructed from multiple sources, which cover diverse scenarios, devices, angles, heights, and ranges. *17-20 are four vehicle clients.
Client Index 0-6 7 8 9 10-11 12-13 14-16 *17-20
Dataset DAIR-V2X [8] LUMPI [9] IPS300+ [10] CAMPUS DAIR-V2X [8]
Scenario Crossroads Cusy Intersection Busy Intersection Campus Roadside Crossroads
Modality Image / PC PC PC PC PC PC Image / PC PC
With GT w w/o w/o w/o w/o w w w/o
@@
Device Jaguar HDL64 Pandar64 PandarQT VLP16 Ruby-Lite Leishen C32 Velodyne 128p
Horizontal FOV (°) 100 360 360 360 360 360 100 100
Vertical FOV (°) 40 26.5 40 104.2 30 40 31 40
Channels (beam) 300 64 64 64 16 80 32 128
Rate (Hz) 10 10 10 10 10 10 10 10
Range (m) 280 120 200 20 100 230 100 245
Range Accuracy (cm) 3 2 2.5 3 3 3 3 3
Installation Height (m) $\sim$6 $\sim$5 $\sim$5 $\sim$5 $\sim$5 5.5 $\sim$6 $\sim$2
Installation Angle(°) 30 0 0 0 0 0 30 0